Part A Two point charges 2.0 cm apart have an electric potential energy - 170 HJ . The total charge is 35 nC. What are the two charges? Express your answers using two significant figures. Enter your answers numerically separated by commas. V AZOO 3 ? 91, 92 = nC
Answer #1

The electric potential enregy 41M2 (-170 * 10-6) * (2 * 10-2) 9 * 109 * q1a2 q1q2 = 37.77 * 10-17 = 37.77 * 10-8 nC